OpenStudy (anonymous):

WILL FAN/MEDAL!!!! A line passes through the point (4,2) and has a slope of 3. Find the equation of the line. A. y=-3-10 B. y= 3x-14** C.y=3x-10 D y=-3x-14 I put a star next to the one I think is correct! Please explain how to do this if I'm not right because I feel like I didn't do this problem correctly!

OpenStudy (anonymous):

Point slope form: \[y-y _{1}=m(x-x _{1})\], where m = slope and (x1, y1) is the point

OpenStudy (anonymous):

y = 2 = 3(x - 4) Solve for y

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

Sorry was supposed to be y - 2 = 3(x - 4)

OpenStudy (anonymous):

So would it be y=-4x+11? I don't know if I did that right or not :( Still new to this >,< @peachpi and @jackmullen55 are you sure? I just got the other answer as the answer to the equation

OpenStudy (anonymous):

I basically took 2 as the y value and 3 as the x value and the m value was -4. but a slope couldn't be negative I thought?! Maybe I just set up the equation wrong?

OpenStudy (anonymous):

A slope can be negative

OpenStudy (anonymous):

Cool hair btw

OpenStudy (anonymous):

Slope can be negative. The point is gives the x and y values (x, y). So x1 = 4, y1 = 2, and The slope is m = 3 y - 2 = 3(x - 4) Distribute 3 y - 2 = 3x - 12 Add 2 to both sides y = 3x + 10

OpenStudy (anonymous):

Thank you so much! I was confused on how to do this but you made it so clear xD so whatever's in the first half before the equal sign is the number we use to add to both sides? and the number after the ='s sign is the number we distribute or does it vary for each equation?! :) Sorry for all the questions I just need to know for future reference!